# Lifelong Learning of Video Diffusion Models From a Single Video Stream

### v2

@@ 4 Model and Training Regime
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 5.5 ±0.2 6.2e-5 ±2e-6 1.63 ±0.03 0.005 ±0.001 5.4 ±0.1 6.2e-5 ±2e-6 1.58 ±0.04 0.005 ±0.0
Lifelong Learning 5.0 ±0.1 6.3e-5 ±1e-6 1.71 ±0.05 0.004 ±0.0 5.0 ±0.2 6.2e-5 ±1e-6 1.68 ±0.01 0.004 ±0.0
Table 1: Lifelong Bouncing Balls (O) performance metrics (lower is better for all) computed across two training and three sampling seeds.
@@
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 6.9 ±0.3 6.5e-5 ±2e-6 1.92 ±0.09 0.007 ±0.002 8.0 ±0.5 6.5e-5 ±2e-6 2.01 ±0.09 0.006 ±0.001
Lifelong Learning 6.1 ±0.3 7.4e-5 ±0.0 1.89 ±0.03 0.005 ±0.0 6.7 ±0.3 7.5e-5 ±1e-6 1.96 ±0.01 0.005 ±0.001
Table 2: Lifelong Bouncing Balls (C) performance metrics (lower is better for all) computed across two training and three sampling seeds.
This setup is compatible with numerous techniques proposed by the continual learning community. To demonstrate this compatibility, we augment the online stream with a minimal experience replay buffer [20] that retains $N_{\text{buffer}}$ past video subsequences chosen through reservoir sampling [77]. $N_{\text{buffer}}$ is typically chosen to be a small fraction of the total data stream size ($\approx 5-20\%$). At training step $t$, the model trains on the $t$-th to $t\!+\!K$-th frames from the data stream as well as $N_{\text{batch}}\!-\!1$ random replay buffer sequences, resulting in the same batch size as standard Offline Learning. We emphasize that replay is a relatively simple and well-established technique that could be substituted with any combination of other continual learning methods. When no replay buffer is used, we fill each training batch with $N_{\text{batch}}$ replications of the $t$-th to $t\!+\!K$-th frames. The diffusion model training procedure samples different values of Gaussian noise for each of these replications, resulting in a reduced-variance estimate of the video diffusion model loss [40] (this can also be done for experience replay to prioritize learning new data).
5 Experiments
This section qualitatively and quantitatively investigates the capabilities of Offline Trained and Lifelong Learned video diffusion models. The learning algorithms use the same batch size and number of gradient steps, equalizing the amount of computation and runtime memory. We refer the readers to Appendix B for additional experiment details.
5.1 Lifelong Bouncing Balls.
As discussed in Section 3, the Bouncing Balls dataset features repetitive and deterministic dynamics, with stationary (Lifelong Bouncing Balls (O)) and non-stationary (Lifelong Bouncing Balls (C)) color dynamics.
@@ 5 Experiments
The models for Lifelong Bouncing Balls (O) and Lifelong Bouncing Balls (C) respectively have 8 and 74 million parameters. The models are trained with a batch size of 2 and are evaluated using 45 frames that they generate conditioned on 5 ground truth frames. Lifelong Learning retains 5 percent of the video stream’s frames in the replay buffer. We measure the sampled videos’ perceptual and temporal coherence using FVD [72] and report the model loss to measure learning progress. We also measure the models’ understanding of the ball movement using minADE [59] and of the color transition using our ColorKL metric (Appendix A), extracting the ball positions and colors from the frames using 2D convolutions.
@@
Method FVD KVD Loss FVD KVD Loss
Offline Learning 93.5 ±1.3 66.6 ±0.9 0.0056 ±0.0 89.2 ±1.0 62.7 ±0.8 0.0056 ±0.0
Lifelong Learning 98.0 ±3.2 62.8 ±2.7 0.0068 ±0.0001 98.8 ±2.4 62.3 ±1.5 0.0059 ±0.0
Table 3: Lifelong 3D Maze performance metrics (lower is better for all) computed across two training and three sampling seeds.
@@
Figure 3: Sample comparison for Lifelong Bouncing Balls (C) where ball colors have changed from red/yellow/green. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(a) and Fig. 3. For Lifelong Bouncing Balls (O), Offline Learning and Lifelong Learning produce models that generate visually compelling ball colors and trajectories, correctly handling different ball-to-ball and ball-to-wall collisions. While no model perfectly recovers the deterministic ground truth trajectories, the videos generated by offline and lifelong learned models are not perceptibly qualitatively different. Because the model’s context size $K=10$ is too small to always condition on two past ball bounces, the models do not perfectly capture the color transition from red to yellow and green. However, the color transition from yellow to red and green to red are captured well, and the models probabilistically transition from red to yellow or green – the best possible modeling choice given the limited context size. Perhaps surprisingly, we observe similar behaviors for Lifelong Bouncing Balls (C). Despite the non-stationary color changes, both learning algorithms generate realistic ball trajectories and colors, regardless of whether we test on frames from the beginning or the end of the video stream.
The quantitative results in Tab. 1 and Tab. 2 mirror the qualitative results. Offline Learning and Lifelong Learning perform comparably on both the stationary Lifelong Bouncing Balls (O) and non-stationary Lifelong Bouncing Balls (C) datasets. Overall, we find that the presence of temporal correlation and the degree of non-repetitiveness do not pose a significant learning challenge to lifelong learning of video diffusion models with Lifelong Learning when the video stream is relatively simple.
@@
Method FVD KVD Loss FVD KVD Loss
Offline Learning 123.2 ±1.1 10.6 ±0.3 0.0330 191.8 ±1.6 14.9 ±0.3 0.0403
Lifelong Learning 107.3 ±1.2 8.3 ±0.2 0.0334 170.7 ±1.4 12.0 ±0.6 0.0408
Table 4: Lifelong PLAICraft performance metrics (lower is better for all) computed across one training and three sampling seeds.
5.2 Lifelong 3D Maze.
The Lifelong 3D Maze dataset – a first-person video stream from a partially observable navigation environment – is a step up in complexity from Lifelong Bouncing Balls. The frames in the video stream are renderings of a 3D space, whereas in Lifelong Bouncing Balls frames consist of 2D objects against a black background. Moreover, as discussed in Sec. 3, the 3D Maze data stream features randomly generated surroundings as well as rare events.
@@
The diffusion models trained in this section have 78 million parameters. The models are trained with a batch size of 4 and are evaluated by generating 40 frames conditioned on 10 ground truth frames. Lifelong Learning retains 5 percent of the video stream’s frames in the replay buffer. We measure the sampled videos’ perceptual and temporal coherence using FVD and KVD [8] using the I3D network [17] from the FVD calculation. We also report the model loss to measure learning progress.
@@
Figure 4: Sample comparison for Lifelong 3D Maze’s rising walls rare event. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(b) and Fig. 4. Both Offline Learning and Lifelong Learning models generate visually compelling maze trajectories, successfully modeling sparsely occurring event sequences such as camera inversion and the rising of the maze walls. While the models occasionally make mistakes such as deforming polyhedral gray rock during camera inversion, these mistakes occur for both models.
Quantitative results appear in Tab. 3. Similar to the qualitative results, the performance metrics are comparable for Offline Learning and Lifelong Learning. Overall, lifelong learned models can capture information on rare events when learning from correlated video frames from a stochastically generated environment, despite heavy data imbalance between rare events and regular maze traversal frames.
5.3 Lifelong PLAICraft.
The Lifelong PLAICraft dataset is the most challenging of the three datasets. Like Lifelong 3D Maze, it is a continual video stream from a partially observable environment. However, the world is more complex with dynamics resulting from interactions between multiple agents.
Setup
The diffusion models trained in this section have 80 million parameters. Lifelong Learning retains 20 percent of the video stream’s frames in the replay buffer. The models are trained with a batch size of 8, of which two elements are the current timestep video frame subsequence, and are evaluated using 10 frames that they generate conditioned on 10 ground truth frames. We measure the sampled videos’ perceptual and temporal coherence using FVD and KVD, and report the model loss to measure learning progress.
Result
Figure 5: Sample comparison for Lifelong PLAICraft’s carrot harvesting sequence. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(c) and Fig. 5. We find that both Offline Learning and Lifelong Learning models capture perceptual details about Minecraft video frames despite having only 80 million parameters and 50 hours of gameplay data to work with. Objects present in every gameplay frame – player name, item bar, and the equipped item – are consistently included in all Offline Learning and Lifelong Learning generations. Compared to the other two datasets, both learning paradigms are less successful at capturing temporal correlations between frames. We hypothesize that a larger model size and careful hyperparameter tuning will bring the sample quality closer to that of the Lifelong Bouncing Balls and Lifelong 3D Mazes models. Nevertheless, the generated video frames from Offline Learning and Lifelong Learning models are not perceptually distinguishable (Appendix G contains additional results).
Quantitative results appear in Tab. 4. On both training and test streams, Offline Learning and Lifelong Learning perform comparably. To summarize, we find that Lifelong Learning can perform comparably to Offline Learning with minimal hyperparameter tuning, even on data streams as complex as Lifelong PLAICraft.
@@ 5.4 Discussion
Given the strong performance of Lifelong Learning, it is natural to wonder if performance could be improved (or worsened) by strengthening (or weakening) the continual learning techniques used as part of the learning paradigm. To that end, in Appendix D we repeat all Lifelong Learning experiments without the use of any replay buffer. Lifelong Learning still obtains competitive performance on repetitive datasets like Lifelong Bouncing Balls (O) and Lifelong 3D Maze. However, there is a more significant performance degradation on the non-repetitive Lifelong Bouncing Balls (C) dataset as well as Lifelong PLAICraft due to catastrophic forgetting.
At the same time, increasing the size of the replay buffer does not result in any significant performance improvement. In Appendix D we see that having access to an unlimited replay buffer – capable of storing all past frames – offers no significant performance improvement over the minimal replay buffer used in this section for all experiments. Taken together, these results suggest that offline performance can be achieved with very little continual learning methodology. We hypothesize that our minimal replay buffer may even be overkill, and simpler regularizations may also yield competitive Lifelong Learned models.
Lastly, to further emphasize that the strong performance of Lifelong Learning is not a byproduct of the generative modeling tasks being unchallenging, we plot the change in the video diffusion models’ test stream quantitative metrics during training for the easiest and hardest datasets, Lifelong Bouncing Balls (O) and Lifelong PLAICraft, in Fig. 6 (similar figures for all datasets are presented in Appendix E). Both Offline and Lifelong Learned models continue to improve performance as training progresses, indicating that the generative modeling tasks are not easy enough to be mastered after the models train on a moderate number of video frames.
(a) Lifelong Bouncing Balls (O) dataset train-time test stream performance metrics.
(b) Lifelong PLAICraft dataset train-time test stream performance metrics.
Figure 6: The evolution of test stream performance metrics during training at different training iterations. The plots show the improvement in model quality as online training progresses. Error bars reflect standard error computed across one training and three sampling seeds.
@@ 6 Related Work
Our work is closely related to continual generative modeling [55, 58, 51, 69, 83, 16, 21], in particular to methods that focus on diffusion models and video generative models. Prior work on lifelong learning of diffusion models [51, 69, 83] focuses on image modeling under the task-based continual learning setup where data grouped into disjoint tasks arrives in large batches. In contrast, our work focuses on lifelong learning of video diffusion models capable of capturing temporal correlations in video frames by learning from a single video stream. Prior work on lifelong learning of video generative models [16, 21] continually learns VAE-based models that can generate future frames again in the task-based continual learning setup. Unlike our setup, these methods assume the presence of rigid task boundaries and the ability to train to convergence on large data batches. In contrast, our lifelong learning datasets do not have notions of tasks, and our models only have access to data as they appear in the stream.
Our work is also related to online learning of sequence processing models [87, 18, 12, 50]. Notably, Carreira et al. [18] learns predictive video models from the concatenation of loosely related videos. In contrast, our work learns generative video models. Liu et al. [50] learns a linear regression-based world model in an online fashion, an approach that we note is not amenable to high-dimensional videos.
Lastly, several recent works have, like us, introduced video datasets for continual learning. Villa et al. [76], Tang et al. [70]’s datasets contain many short videos that can be learned under a task-based continual learning setup. Carreira et al. [18]’s datasets construct one very long data stream by concatenating multiple short-to-medium length videos, but their data are not publicly available. Singh et al. [68]’s dataset, while not originally developed for continual learning, contains a series of short-to-medium-length videos of the real world that were collected via Google Glass. We note that their dataset has many semantic discontinuities, whereas our datasets come from a continuous stream under a single environment.

### v3

@@ 5.1 Lifelong Bouncing Balls
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Offline Learning (U) 4.5 ±0.1 6.3 ±0.1 1.74 ±0.03 0.006 ±0.001 4.7 ±0.2 6.2 ±0.2 1.71 ±0.06 0.006 ±0.001
Lifelong Learning (U) 4.9 ±0.2 6.3 ±0.2 1.82 ±0.03 0.005 ±0.000 4.7 ±0.2 6.2 ±0.3 1.81 ±0.03 0.005 ±0.000
Offline Learning (T) 5.4 ±0.2 6.8 ±0.3 2.25 ±0.12 0.006 ±0.001 5.1 ±0.2 6.7 ±0.3 2.16 ±0.12 0.006 ±0.001
Lifelong Learning (T) 5.5 ±0.1 6.6 ±0.2 2.22 ±0.05 0.005 ±0.001 5.0 ±0.3 6.5 ±0.2 2.16 ±0.06 0.005 ±0.001
Table 1: Lifelong Bouncing Balls (O) metrics computed across two training and three sampling seeds. The configurations (U) and (T) denote U-Net and Transformer-based diffusion architecture.
@@
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Offline Learning (U) 5.8 ±0.3 6.5 ±0.1 2.04 ±0.09 0.007 ±0.002 5.9 ±0.2 6.5 ±0.1 2.14 ±0.10 0.007 ±0.001
Lifelong Learning (U) 5.0 ±0.1 7.4 ±0.1 2.03 ±0.00 0.005 ±0.001 5.7 ±0.2 7.5 ±0.1 2.06 ±0.00 0.005 ±0.000
Offline Learning (T) 5.0 ±0.2 8.4 ±0.1 2.24 ±0.05 0.006 ±0.000 5.6 ±0.1 8.5 ±0.1 2.19 ±0.02 0.006 ±0.000
Lifelong Learning (T) 5.3 ±0.1 8.1 ±0.0 2.29 ±0.03 0.006 ±0.001 6.4 ±0.3 8.1 ±0.0 2.27 ±0.03 0.006 ±0.000
Table 2: Lifelong Bouncing Balls (C) metrics computed across two training and three sampling seeds. The configurations (U) and (T) denote U-Net and Transformer-based diffusion architecture.
Result
Qualitative results appear in fig. 1(a) and fig. 3(a). For Lifelong Bouncing Balls (O), Offline Learning and Lifelong Learning produce models that generate visually compelling ball colors and trajectories, correctly handling different ball-to-ball and ball-to-wall collisions. While no model perfectly recovers the deterministic ground truth trajectories, the videos generated by offline and lifelong learned models are not perceptibly different in quality. Because the model’s context size $K=10$ is too small to always condition on two past ball bounces, the models do not perfectly capture the color transition from red to yellow and green but instead probabilistically transition – the best possible modeling choice given the limited context size. We observe the same behaviors for Lifelong Bouncing Balls (C). Despite the non-stationary color changes, both learning algorithms generate realistic ball trajectories and colors, regardless of whether we test on frames from the beginning or the end of the video stream. The quantitative results in table 1 and table 2 mirror the qualitative results. Offline Learning and Lifelong Learning perform similarly on both the stationary Lifelong Bouncing Balls (O) and non-stationary Lifelong Bouncing Balls (C) datasets. Overall, the presence of temporal correlation and the degree of non-repetitiveness does not pose a significant learning challenge to Lifelong Learning when the video stream is relatively simple.
5.2 Lifelong 3D Maze
The Lifelong 3D Maze dataset is a step up in complexity from Lifelong Bouncing Balls. The video stream frames are renderings of a 3D space and feature stochastic dynamics and rare events.
@@ 5.2 Lifelong 3D Maze
The U-Net and Transformer-based diffusion models from this section respectively have 78 million and 110 million parameters. All models are trained with a batch size of 4, of which two elements are the current timestep video frame subsequence, and are evaluated using 40 frames autoregressively generated from conditioning on 10 ground truth frames. Lifelong Learning retains 5 percent of the video stream’s frames in the replay buffer. We measure perceptual and temporal coherence using FVD and JEDi [53] and learning progress using the diffusion loss.
Result
Qualitative results appear in fig. 1(b) and fig. 3(b). Both Offline Learning and Lifelong Learning models generate coherent maze trajectories, successfully modeling sparsely occurring event sequences such as camera inversion and the rising of the maze walls. Both models make the occasional mistake, such as deforming polyhedral gray rocks during camera inversion. Quantitative results appear in table 3. We find that the Offline and Lifelong Learning metrics are generally similar, and neither methods significantly outperform the other across all metrics, architectures, and stream types. Overall, we see on qualitative and quantitative fronts that Lifelong Learning performs well and can capture information on rare events when learning from correlated frames from a stochastically generated environment, despite heavy data imbalance between rare events and regular maze traversal frames.
@@
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning (U-Net) 34.2 ±1.3 0.083 ±0.002 0.005 ±0.0 28.5 ±1.2 0.060 ±0.001 0.005 ±0.0
Lifelong Learning (U-Net) 41.2 ±0.4 0.073 ±0.006 0.006 ±0.0 30.7 ±0.1 0.061 ±0.003 0.006 ±0.0
Offline Learning (Transformer) 30.2 ±0.9 0.088 ±0.005 0.006 ±0.0 32.6 ±1.0 0.082 ±0.005 0.006 ±0.0
Lifelong Learning (Transformer) 32.1 ±0.9 0.082 ±0.005 0.006 ±0.0 30.2 ±0.8 0.070 ±0.001 0.006 ±0.0
Table 3: Lifelong 3D Maze metrics computed across one training and three sampling seeds.
Train Stream Test Stream
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning (U-Net) 15.3 ±0.2 0.071 ±0.000 0.029 ±0.0 17.9 ±0.3 0.102 ±0.001 0.026 ±0.0
Lifelong Learning (U-Net) 17.0 ±0.2 0.072 ±0.000 0.030 ±0.0 21.8 ±0.3 0.106 ±0.000 0.027 ±0.0
Offline Learning (Transformer) 10.4 ±0.2 0.023 ±0.000 0.027 ±0.0 11.2 ±0.1 0.036 ±0.000 0.024 ±0.0
Lifelong Learning (Transformer) 10.7 ±0.1 0.023 ±0.000 0.029 ±0.0 12.9 ±0.1 0.039 ±0.000 0.025 ±0.0
Table 4: Lifelong Drive metrics computed across one training and three sampling seeds.
5.3 Lifelong Drive
The Lifelong Drive dataset features real-world perceptual data captured in a 3D environment similar to Lifelong 3D Maze, but with higher frame resolution and greater perceptual complexity.
Setup
The U-Net and Transformer-based diffusion models in this section respectively have 80 and 300 million parameters. Lifelong Learning retains 20 percent of the video stream’s frames in the replay buffer. All models are trained with a batch size of 8, of which two elements are the current timestep video frame subsequence, and evaluation is performed on videos with 20 frames, 10 of which are model-generated given the preceding 10 frames. We report FVD, JEDi, and model loss.
@@ 5.3 Lifelong Drive
See fig. 1(c) and fig. 3(c) for qualitative results. Similar to 3D Maze, Lifelong Learning models produce realistic dynamics of moving in three dimensions. The generated videos from models trained with either algorithm are realistic and are capable of modeling actions like lane changes, though there exists a mild autoregressive drift less present in simpler datasets (ex. fig. 3(c) bridge shape). Quantitative results appear in table 4. Offline Learning and Lifelong Learning perform similarly across metrics, architectures, and stream types. We see that changing the model architecture and size has a much more significant impact on model performance than the learning algorithm, suggesting that video diffusion models can be effectively lifelong learned from an autocorrelated video stream.
5.4 Lifelong PLAICraft
@@
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning (U-Net) 59.7 ±0.2 0.874 ±0.007 0.034 ±0.0 119.8 ±1.3 1.119 ±0.004 0.042 ±0.0
Lifelong Learning (U-Net) 62.9 ±0.6 0.876 ±0.003 0.034 ±0.0 130.8 ±1.1 1.189 ±0.003 0.042 ±0.0
Offline Learning (Transformer) 23.7 ±0.1 0.209 ±0.001 0.024 ±0.0 27.7 ±0.3 0.223 ±0.002 0.024 ±0.0
Lifelong Learning (Transformer) 21.9 ±0.4 0.197 ±0.003 0.025 ±0.0 26.3 ±0.3 0.208 ±0.001 0.024 ±0.0
Table 5: Lifelong PLAICraft metrics computed across one training and three sampling seeds.
The Lifelong PLAICraft dataset is the most challenging in terms of temporal dynamics out of the four datasets. The video stream is continuously recorded from a partially observable environment containing multi-agent interaction and a highly diverse perceptual and action space.
@@
The U-Net and Transformer-based diffusion models in this section respectively have 80 and 300 million parameters. Lifelong Learning retains 20 percent of the video stream’s frames in the replay buffer. All models are trained with a batch size of 8, of which two elements are the current timestep video frame subsequence, and evaluation is performed on videos with 10 frames, 5 of which are model-generated conditioned on the preceding 5 frames. We report FVD, JEDi, and model loss.
@@
See fig. 1(d) and fig. 3(d) for qualitative results. Both Offline Learning and Lifelong Learning models capture perceptual details about Minecraft video frames. Objects present in every gameplay frame (player name, item bar, and the equipped item) are consistently included in all Offline and Lifelong generations, and even chat logs are faithfully rendered (fig. 1(d)). Frequently observed temporal dynamics such as item movement during walking and flashing of the bottom right recording dot are captured better than player movement and actions. We hypothesize that a larger model size and careful hyperparameter tuning achievable on commercial-scale compute will significantly improve the sample quality for both learning algorithms. Nevertheless, the generated videos from offline and lifelong learned models are not perceptually distinguishable in quality. Quantitative results appear in table 5. On both training and test streams, Offline Learning and Lifelong Learning perform similarly, with neither approach consistently outperforming the other across metrics and architectures. In summary, we find that Lifelong Learning can perform comparably to Offline Learning with minimal hyperparameter tuning, even on videos as complex as Lifelong PLAICraft.
@@ 5.5 Discussion
The reported metrics for Offline and Lifelong Learning across all datasets are remarkably similar. While the confidence intervals for both learning algorithms do not always overlap for every metric, dataset, and model architecture configuration, the performance metric means are largely similar and other control variables such as model architecture, parameter size, and dataset have substantially stronger effects on the metrics. Combined with the qualitative indistinguishability of offline versus lifelong model samples, we conclude that video diffusion models can be effectively lifelong-learned.
Ablation studies.
We investigate model sensitivity to the replay buffer size for every dataset in Appendix D. We find that storing 5 to 20 percent of video stream frames is often sufficient and that unlimited replay buffer size does not lead to significantly stronger results for our diverse group of datasets. We also find that not having a replay buffer impedes model performance and that this behavior is more pronounced when the training stream is nonstationary. Furthermore, we plot the fully trained U-Net models’ performance metrics on videos from different parts of the training stream for all datasets in Appendix G. We observe that the addition of the replay buffer causes Lifelong Learning to perform similarly to Offline Learning on video subsequences from different training stream timesteps, suggesting that forgetting is not a major obstacle given the replay buffer. These results suggest that experience replay with a modest buffer size is a strong and simple baseline when lifelong learning video diffusion models are in memory-constrained settings.
In addition, inspired by online continual learning methods that combine replay with stabilizing optimizers [86, 75], we evaluate a lifelong learning approach that minimizes eq. 2’s replay loss with the Orthogonal AdamW optimizer [34] instead of AdamW and report the results in Appendix E. The resulting models perform comparably to regular experience replay, suggesting that our findings are not highly sensitive to the specifics of optimizer configuration.
Does learning occur throughout the entire data stream?
To demonstrate that learning realistic models on these datasets is non-trivial, we plot the change in the video diffusion models’ test stream quantitative metrics during training for all datasets in Appendix F. Both Offline and Lifelong Learning continue to improve in performance as training progresses, indicating that the generative modeling tasks cannot easily be mastered after the models train on a moderate number of video frames.
@@ 6 Related Work
Our work is closely related to continual generative modeling [58, 60, 54, 71, 87, 17, 21], in particular to methods that focus on diffusion models and video generative models. Prior work on diffusion model lifelong learning [54, 71, 87, 50] focuses on image modeling under the task-based continual learning setup where data grouped into disjoint tasks arrives in large batches. In contrast, our work focuses on lifelong learning of video diffusion models capable of capturing temporal correlations in video frames by learning from a single video stream. Prior work on video generative model lifelong learning [17, 21] trains VAE models that can generate future frames again in the task-based continual learning setup. Unlike our setup, these methods assume the presence of rigid task boundaries and the ability to train to convergence on large data batches. In contrast, our lifelong learning datasets do not have notions of tasks, and our models only have access to data as they appear in the stream.
Our work is also related to online learning of sequence processing models [91, 18, 13, 51, 34]. Notably, Carreira et al. [18], Han et al. [34] learns discriminative video prediction models from the concatenation of loosely related videos. In contrast, our work learns generative video models from a single autocorrelated stream. Liu et al. [51] learns a linear regression-based world model in an online fashion, an approach that we note is not amenable to high-dimensional videos.
Lastly, prior work has, like us, introduced video datasets for continual learning. Villa et al. [80], Tang et al. [72]’s datasets contain many short videos that can be learned under a task-based continual learning setup. Carreira et al. [18]’s datasets construct one very long data stream by concatenating multiple short-to-medium length videos, but their data are not publicly available. Singh et al. [70]’s dataset, while not originally developed for continual learning, contains a series of short-to-medium-length videos of the real world that were collected via Google Glass. We note that their dataset has semantic discontinuities, whereas our datasets come from single, continuous video streams.